# CI/CD Configuration Practices in Open-Source Android Apps: An Empirical Study

### v1

@@ 3.1.1. Collecting Android projects from F-Droid
The F-Droid (F-Droid, [n.d.]) store serves as a prominent repository for Android applications, offering a diverse selection of mobile applications. GitHub projects were previously studied in prior work as GitHub offers extensive data for analyzing developer activities, such as the commit messages and reported issues. Hence, we filter applications that contain links to their GitHub project. Following this filtration, we identify $3,911$ mobile apps.
@@ 3.1.2. Collecting Android projects from GitHub
To gather data from GitHub, we utilize the official GitHub API [^5] https://docs.github.com/en/rest. Given the enormous base of GitHub projects, we focus specifically on projects that are classified under the ‘android’ topic [^6] https://github.com/topics/android, similar to Liu et al. (2020). This step, conducted in May 2023, provides us with $105K$ GitHub projects. To further refine our dataset, we exclude projects without a user interface by verifying the presence of an AndroidManifest.xml file with an activity tag. This filtration step ensures that the selected projects are indeed Android applications, thereby enhancing the accuracy and relevance of our dataset for analysis. At the end of this step, we identify $40,886$ Android-related projects.
@@ 3.1.3. Merging all projects from the two sources
After collecting Android projects from both GitHub and the F-Droid store, we remove duplicated projects. As such, we identify $43,283$ unique projects.
@@ 3.1.5. Removing toy projects
Android projects containing YML files do not necessarily represent an actual mobile application. For example, tutorial and library projects include example apps developed for demonstration purposes that have identical structure to actual mobile apps, including AndroidManifest.xml and YML files, but they are not maintained or updated further. We conduct a semi-automated approach combining manual and automated analysis to identify and exclude toy projects from our dataset. First, through collaborative discussion sessions, two co-authors manually examined a statistically significant sample of 93 projects (a confidence level of $95\%$ percent and a confidence interval of ±$\pm 10$), checking their descriptions and AndroidManifest.xml paths to identify an initial set of keywords indicative of toy projects (e.g., ‘example’, ‘sample’, ‘library’, and ‘tutorial’). Then, based on the findings from our manual analysis, we develop heuristics in the form of a Python script to automate the identification of toy projects within our dataset. These scripts utilized the keywords identified during the manual process to filter projects programmatically. Following the automated filtering process, the two co-authors independently performed a manual validation of another statistically significant sample of projects (an additional 93 projects) to confirm the effectiveness and accuracy of the automated filtering. This validation step ensured that our automated approach reliably distinguished between toy projects and real mobile apps in our dataset. We use two codes in which the raters indicate whether there is a match or mismatch between the automatically generated and manually assigned labels. We discuss the cases in which there is a disagreement or mismatch between manual and automated filtering. The first few iterations show $14\%$-$19\%$ mismatch and disagreement cases. We examine those cases and resolve disagreements using consensus meetings, refine our heuristics, and then repeat the same process for the remaining unfiltered projects. Eventually, we achieve a strong degree of observed agreement ($93\%$) in manually validating the automatically filtered projects.
@@
After removing toy projects, we identify $2,564$ projects that adopt CI/CD practices. Table 2 shows our final set of projects and CI/CD services. We exclude CI/CD services (e.g., AppVeyor, Azure Pipelines, and Bitbucket) with fewer than $1.00\%$ projects, as this percentage is not representative enough to draw conclusions. Finally, we obtain $2,719$ Android apps adopting four different CI/CD services, namely GitHub Actions [^7] https://github.com/features/actions, Travis CI [^8] https://www.travis-ci.com, CircleCI [^9] https://circleci.com, and GitLab CI/CD [^10] https://docs.gitlab.com/ee/ci. The majority (about $90\%$) of these projects are developed using the Kotlin, Java, and Dart programming languages. Table 3 shows the full list of programming languages with their corresponding number of projects across CI/CD services.
@@ 3.1.6. Crawl all CI/CD configuration files
For the final set of projects, we crawl the configuration YML files associated with the adopted CI/CD services. We collect a total of $4,338$ files. We should note that (a) there are projects that adopt multiple services at the same time and (b) other projects have multiple YML files for one service. This led to having more configuration files compared to the number of projects (an average of 1.7 files per project).
@@
GitHub Actions 1,685 (58.52%) 3,268 (75.33%)
Travis CI 749 (28.42%) 749 (17.27%)
CircleCI 197 (9.56%) 197 (4.54%)
GitLab CI/CD 88 (2.34%) 88 (2.03%)
Table 3. Programming languages with their corresponding number of projects across CI/CD services
@@
Kotlin 770 186 107 24 1,092
Java 477 455 62 44 1,044
Dart 228 20 4 3 257
C++ 46 12 1 7 77
TypeScript 48 9 1 1 61
JavaScript 36 18 4 1 60
C 21 15 1 4 46
@@ 3.2.1. Extracting CI/CD configurations
We extract all relevant YML files representing CI/CD configurations from projects that adopt CI/CD practices. Subsequently, we process these files using the PyYAML package [^11] https://pypi.org/project/PyYAML in Python. Each file contains pairs of keys (referred to as directives) and values. Directives like ‘name:’, ‘run:’, and others are the keys that define specific instructions or configurations in a CI/CD workflow. Each directive has a corresponding value that specifies how the workflow should operate, such as the name of a job, a command to run, or other settings. Through parsing each of these files, we extract the directives and their corresponding values by combining each directive with its precedent directives, thus representing their full hierarchical path within the configuration. For example, a value, such as “./gradlew dependencies”, associated with a directive like ‘jobs.build.steps.run’ instead of ‘run’ alone, resulting in a complete path that distinguishes directives across the CI/CD process. This approach helps capture comprehensive information on the CI/CD setups implemented across various projects, making it easier to understand the context and connections of each directive within the overall configuration. In total, our dataset contains a set of $24,976$ unique pairs of directives/values.
@@ 3.2.2. Categorizing configuration directives
    . Configuration directives involved in the build process and packaging of the Android application. This typically includes compiling the source code and generating build artifacts. Listing 3 gives an example of directives under this category in GitHub Actions and Travis CI.
@@ 4. Results
4.1. RQ1: How do CI/CD configurations vary in terms of complexity and tool usage?
@@
CI/CD practices have become integral to modern software development, facilitating rapid and reliable delivery of software products. However, as CI/CD pipelines grow in complexity and diversity, understanding the complexities of configuration practices and tool usage across different phases of these pipelines becomes critical for several reasons. Firstly, it allows organizations to optimize their CI/CD workflows by identifying potential challenges or inefficiencies in configuration practices. Secondly, it provides insights into the prevalent tools and technologies used by practitioners, enabling informed decision-making regarding tool adoption and integration. Moreover, by shedding light on the disparities in configuration complexity and tool usage, this RQ contributes to the broader discourse on CI/CD best practices and service comparisons. It provides stakeholders, including software developers, DevOps engineers, and project managers, with actionable insights to enhance the efficiency and effectiveness of their CI/CD processes.
@@
To address this RQ, we conduct a thorough analysis of CI configurations, focusing on the complexity across different services and categories. We also identify the common tools developers use to test and deploy their mobile apps.
Analyzing CI/CD configuration complexity. We investigate the complexity of configuration files among various CI/CD services to recognize differences. Our analysis primarily focuses on providing descriptive statistics, such as median, maximum value, minimum value, and other relevant measures. Moreover, to identify significant differences between the complexities of CI/CD configurations of different CI/CD services, we utilize the Wilcoxon test. This non-parametric test is particularly suited for comparing two paired groups, which aligns well with our objective of evaluating differences among the services. Finally, we analyze the complexity of directives under different categories to understand which categories require more configuration and attention across various services. To do this, we calculate the number of directive/value pairs under each category for each project independently. Given that configuration size varies across projects, we calculate the percentage of each category for each project. We use boxplots to visualize these percentages, highlighting the descriptive statistics overall and for each CI/CD service.
@@
Observation 1.1: The complexity of configuration files vary across CI/CD services, with a median of 19 configuration directives per file, as we observe that GitLab CI/CD has the highest value while Travis CI service has the lowest complexity level. Our analysis of configuration directives uncovers significant differences in file complexities among CI/CD services, indicating variations in the utilization of configuration directives across these services. The median number of directives per file is $19$, with distinctive differences observed across CI/CD services. Specifically, GitLab CI/CD stands out with a higher directive count of $25.5$, whereas Travis CI shows a lower count of $16$. Other services fall in between, with GitHub Actions at $20$, and CircleCI at $18$.
Our results reveal that, overall, all CI/CD services have high variability in complexity with many outliers indicating some configuration files are significantly more complex. This implies that those services offer a lot of flexibility and functionality for configuring CI/CD pipelines, but also require more attention and expertise to manage the complexity. Some configuration files may be unnecessarily complex or inefficient, and could benefit from simplification or optimization.
Our investigation reveals a significant difference (p-value $<0.0001$) in the complexities of CI/CD configuration files across various programming languages, indicating that at least one programming language exhibits an observable difference in complexity compared to others. Furthermore, we observe a significant difference (p-value $<0.0001$) in the complexity of files across CI/CD services. Therefore, developers should make informed decisions about which CI/CD service to adopt, carefully considering the balance between the required features and the potential complexity of initializing and maintaining the configuration files.
@@
[…]
@@
Observation 1.3: Nearly half of the projects implement automated testing processes, with the predominant standard testing tool, through Gradle, accounting for about 78% of the total. Among the analyzed projects, 1,223 (47.7%) use automated testing processes as part of the CI/CD pipeline. Table 4 shows the top 10 testing tools used by the projects in our dataset. We should note that multiple testing tools can be used in a single project, resulting in cumulative percentages that exceed 100%. The results show a preference for adopting testing practices through standard tools like Gradle in Android projects. The results are consistent across CI/CD services. This trend likely arises from the ease of integration and familiarity, leading to evidently more efficient testing workflows. We observe that unit testing is by far the most commonly used practice, with 78.3% of projects incorporating it. This highlights the importance of unit testing in software development, as it enables developers to verify code correctness at a granular level. This suggests that unit testing is deemed essential for maintaining code reliability and is likely prioritized due to its effectiveness in identifying defects early and its seamless integration with continuous integration CI/CD pipelines.
Android-specific tools, such as Android Lint (28.7%), testing on connected devices (14.5%), and the Android Emulator (9.6%), show a relatively lower adoption compared to unit testing, despite their role in ensuring platform compliance and optimizing performance. The greater reliance on Android Lint suggests an interest of Android developers in performing static code analysis to detect issues before runtime. However, the low adoption rates of connected devices and emulators might indicate that many teams prefer more lightweight and automated approaches, possibly relying on simulated environments or static checks before moving to next stages.
@@
Finally, tools like Firebase (5.1%), Fastlane (2.2%), and SonarQube (2.1%) show even lower adoption rates, despite their popularity as third-party testing platforms. Firebase’s usage likely reflects a focus on cloud-based testing and monitoring, particularly for mobile applications, but its usage remains relatively limited in our datasets. Fastlane and SonarQube, which facilitate CI/CD automation and comprehensive code quality analysis respectively, appear to be more specialized tools that are leveraged in specific cases, possibly due to their integration complexity.
@@
Testing Tool Configuration Examples # (%)

### v2

@@ 3.1.1. Collecting Android projects from F-Droid
The F-Droid (F-Droid, [n.d.]) store serves as a prominent repository for Android apps, offering a diverse selection of mobile apps. GitHub projects were previously studied in prior work as GitHub offers extensive data for analyzing developer activities, such as the commit messages and reported issues. Hence, we select apps that contain links to their GitHub project. Following this filtration, we identify $3,911$ mobile apps.
@@ 3.1.2. Collecting Android projects from GitHub
To gather data from GitHub, we utilize the official GitHub API [^13] https://docs.github.com/en/rest. Given the enormous base of GitHub projects, we focus specifically on projects that are classified under the ‘android’ topic [^14] https://github.com/topics/android, similar to Liu et al. (2020). This step, conducted in May 2023, provides us with $108k$ GitHub projects. To further refine our dataset, we exclude projects without a user interface by verifying the presence of an AndroidManifest.xml file with an activity tag. This filtration step ensures that the selected projects are indeed Android apps, thereby enhancing the accuracy and relevance of our dataset for analysis. At the end of this step, we identify $40,886$ Android-related projects.
@@ 3.1.3. Merging all projects from the two sources
After collecting Android projects from both GitHub and the F-Droid store, we remove duplicate projects and those that have zero GitHub stars, and identify $43,283$ unique projects.
@@ 3.1.5. Removing toy projects
Android projects containing YML files do not necessarily represent an actual mobile app. For example, tutorial and library projects include example apps developed for demonstration purposes that have a structure identical to actual mobile apps, including AndroidManifest.xml and YML files, but they are not maintained or updated further. We perform manual and automated analysis to identify and exclude toy projects from our dataset. First, through collaborative discussion sessions, two raters (co-authors of this paper with practical experience in software engineering and Android app development) manually examined a statistically significant random sample of 93 projects (a confidence level of $95\%$ and a confidence interval of ±$\pm 10$), reviewing their repository descriptions and AndroidManifest.xml paths to identify an initial set of indicative keywords (e.g., ‘example’, ‘sample’, ‘library’, and ‘tutorial’). Based on this analysis, we develop heuristic rules in the form of a Python script to automate the detection of toy projects within our dataset using these keywords.
To assess the effectiveness of our automated filtering, the same two co-authors independently validate another statistically significant random sample of 93 projects. Each rater indicates whether there is a match or mismatch between the automated label and their manual judgment. Disagreements or mismatches were discussed in consensus meetings, with input from the third co-author when needed. In these two initial iterations, we observed $14\%$-$19\%$ mismatch rates between the manual and automated labeling, prompting refinement of our heuristics. This process of validation, refinement, and re-evaluation was repeated iteratively six times for the remaining unfiltered projects until a $93\%$ accuracy rate was achieved in manually validating the automatically filtered projects, ensuring that our final dataset reliably excluded toy projects and retained real Android mobile apps. After removing toy projects, we identify $2,719$ projects that adopt CI/CD practices.
@@
We exclude CI/CD services (e.g., AppVeyor, Azure Pipelines, and Bitbucket) with fewer than $1.00\%$ projects, as this percentage is not representative enough to draw conclusions. Finally, we obtain $2,557$ Android apps adopting four different CI/CD services, namely GitHub Actions [^15] https://github.com/features/actions, Travis CI [^16] https://www.travis-ci.com, CircleCI [^17] https://circleci.com, and GitLab CI/CD [^18] https://docs.gitlab.com/ee/ci. The majority (about $90\%$) of these projects are developed using the Kotlin, Java, and Dart programming languages. Table 3 shows the full list of programming languages with their corresponding number of projects across CI/CD services.
Table 2 shows our final set of projects and CI/CD services. To provide an overview of our dataset, we present boxplots showing the distribution of key repository-level metrics for our studied projects in Figure 3. Each boxplot summarizes one metric, such as the codebase size (in megabytes), number of commits, pull requests, issues, and popularity indicators like stars and forks. This visualization highlights the large variance across projects (with medians highlighted in orange), particularly in contribution and popularity metrics, with some repositories exhibiting orders-of-magnitude differences.
Figure 3. Boxplots showing log-scaled distributions of repository-level metrics the $2,557$ studied projects, with medians annotated and highlighted in orange color
@@ 3.1.6. Crawl all CI/CD configuration files
For the final set of projects, we crawl the configuration YML files associated with the adopted CI/CD services. We collect a total of $4,295$ files. We should note that (a) there are projects that adopt multiple services at the same time, and (b) other projects have multiple YML files for one service. This led to having more configuration files compared to the number of projects (an average of 1.7 files per project).
@@
GitHub Actions 1,670 (65%) 3,263 (76%)
Travis CI 755 (30%) 748 (17%)
CircleCI 198 (8%) 196 (5%)
GitLab CI/CD 87 (3%) 88 (2%)
Total 2,557 4,295
Table 3. Top 10 programming languages in our dataset with their corresponding number of projects across CI/CD services
@@
Kotlin 770 188 109 24 1,091
Java 477 461 62 44 1,044
Dart 231 21 4 3 259
C++ 46 12 2 7 67
JavaScript 36 18 4 1 59
TypeScript 48 9 1 1 59
C 21 16 1 4 42
@@ 3.2.1. Extracting CI/CD configurations
We extract all relevant YML files representing CI/CD configurations from projects that adopt CI/CD practices. Subsequently, we process these files using the PyYAML package [^19] https://pypi.org/project/PyYAML in Python. Each file contains pairs of keys (referred to as directives) and values. Directives like ‘name:’, ‘run:’, and others are the keys that define specific instructions or configurations in a CI/CD workflow. Each directive has a corresponding value that specifies how the workflow should operate, such as the name of a job, a command to run, or other settings. Through parsing each of these files, we extract the directives and their corresponding values by combining each directive with its preceding directives, thus representing their full hierarchical path within the configuration. For example, a value, such as “./gradlew dependencies”, associated with a directive like ‘jobs.build.steps.run’ instead of ‘run’ alone, resulting in a complete path that distinguishes directives across the CI/CD process. This approach helps capture comprehensive information on the CI/CD setups implemented across various projects, making it easier to understand the context and connections of each directive within the overall configuration. In total, our dataset contains a set of $24,976$ unique pairs of directives/values.
@@ 3.2.2. Categorizing configuration directives
    . Configuration directives involved in the build process and packaging of an Android app. This typically includes compiling the source code and generating build artifacts. Listing 3 gives an example of directives under this category in GitHub Actions and Travis CI.
@@
3.3. Comparisons with CI/CD Configuration Practices in General
We conduct a thorough review of findings reported in prior studies on CI/CD configuration practices across both open-source and industrial projects, regardless of application domain. To do this, we begin with a recent systematic literature review (SLR) on build optimization (Aïdasso et al., 2025) as a base reference, and then perform further cross-referencing to identify studies that specifically address CI/CD configuration. This approach enables us to establish meaningful comparisons with our findings on Android apps and to contextualize the patterns observed in Android-specific pipelines against broader trends and challenges. By examining prior research on configuration smells, anti-patterns, and service-specific practices, we aim for our analysis to capture both domain-specific characteristics and generalizable insights into CI/CD configuration.
@@ 4. Results
4.1. RQ1: How do CI/CD configurations vary in complexity and tool usage for Android apps?
@@
CI/CD practices have become integral to modern software development, facilitating rapid and reliable delivery of software products. However, as CI/CD pipelines grow in complexity and diversity, understanding the complexities of configuration practices and tool usage across different phases of these pipelines becomes critical for several reasons. Firstly, it allows organizations to optimize their CI/CD workflows by identifying potential challenges or inefficiencies in configuration practices. Secondly, it provides insights into the prevalent tools and technologies used by practitioners, enabling informed decision-making regarding tool adoption and integration. Moreover, by shedding light on the disparities in configuration complexity and tool usage, this RQ contributes to the broader discourse on CI/CD best practices and service comparisons. It provides stakeholders, including software developers, CI/CD engineers, and project managers, with actionable insights to enhance the efficiency and effectiveness of their CI/CD processes.
@@
To address this RQ, we conduct a thorough analysis of CI/CD configurations, focusing on the complexity across different services and categories. We also identify the common tools developers use to test and deploy their mobile apps.
Analyzing CI/CD configuration complexity. We investigate the complexity of configuration files among various CI/CD services to recognize differences. Our analysis primarily focuses on providing descriptive statistics, such as median, maximum value, minimum value, and other relevant measures. Moreover, to identify significant differences between the complexities of CI/CD configurations of different CI/CD services, we utilize the Wilcoxon test (Wilks, 2011). This non-parametric test is particularly suited for comparing two paired groups, which aligns well with our objective of evaluating differences among the services. We also use the Kruskal–Wallis test (Tomczak and Tomczak, 2014) to compare groups, and estimate the effect size using eta squared ($\eta^{2}$) and report the Kruskal–Wallis statistic ($H$), following recommendations for non-parametric data. This provides an interpretable measure of the proportion of variance explained by the grouping variable. Finally, we analyze the complexity of directives under different categories to understand which categories require more configuration and attention across various services. To do this, we calculate the number of directive/value pairs under each category for each project independently. Given that configuration size varies across projects, we calculate the percentage of each category for each project. We use boxplots to visualize these percentages, highlighting the descriptive statistics overall and for each CI/CD service.
@@
Observation 1.1: CI/CD configuration is not standardized, and its complexity varies across CI/CD services, with a median of 19 configuration directives per file, as we observe that GitLab CI/CD has the highest value while Travis CI service has the lowest complexity level. Our analysis of configuration directives uncovers significant differences in file complexities among CI/CD services, indicating variations in the utilization of configuration directives across these services. The median number of directives per file is $19$, with distinctive differences observed across CI/CD services. Specifically, GitLab CI/CD stands out with a higher directive count of $25.5$, whereas Travis CI shows a lower count of $16$. Other services fall in between, with GitHub Actions at $20$, and CircleCI at $18$.
Our results reveal that, overall, all CI/CD services have high variability in complexity, with many outliers indicating some configuration files are significantly more complex. This implies that those services offer a lot of flexibility and functionality for configuring CI/CD pipelines, but they also require more attention and expertise to manage the complexity. Some configuration files may be unnecessarily complex or inefficient, and could benefit from simplification or optimization.
Our results reveal a significant difference (p-value ¡ 0.0001) in the complexities of CI/CD configuration files across various programming languages ($H$ = 429.06, $\eta^{2}$ = 0.083), indicating a moderate effect size and suggesting that at least one programming language exhibits a substantial difference in complexity compared to others. Furthermore, we observe a significant difference (p-value ¡ 0.0001) in the complexity of files across CI/CD services ($H$ = 160.87, $\eta^{2}$ = 0.032), with a small to moderate effect size, highlighting meaningful variations in complexity between different CI/CD services. Therefore, developers should make informed decisions about which CI/CD service to adopt, carefully considering the balance between the required features and the potential complexity of initializing and maintaining the configuration files.
Compare & Contrast with Prior Research Findings (1.1): Overall, our results align with prior research indicating that CI configurations were deemed complex and generally challenging for developers, not only in Android apps (Rostami Mazrae et al., 2023; Widder et al., 2019; Hilton et al., 2017; Kavaler et al., 2019; Zhang et al., 2018; Saroar and Nayebi, 2023). For example, Zhang et al. (2018) indicated that developers can become frustrated with specific CI/CD services and discouraged from adopting CI practices, even when more customized options are available. Some studies, based on developer feedback, linked such complexity to YML ’s syntax for being confusing and not easy to write configurations with it (Saroar and Nayebi, 2023). The lack of standardization across different CI/CD services was also reported (Rostami Mazrae et al., 2023), which made migration between services difficult and error-prone, as developers must navigate differing semantics and feature implementations. There is a learning curve in understanding all the configuration options, and it can be complex to configure even a simple CI pipeline (Hilton et al., 2017). In the context of our study, Android apps can often require more complex build setups, platform-specific dependencies, and integration with mobile-specific tools (e.g., emulators), making their CI/CD configuration worth investigating across various services.
@@
Compare & Contrast with Prior Research Findings (1.2): Prior research indicated that the primary focus of CI configuration in software projects is on environments and containers setup (Zhang et al., 2022; Gallaba and McIntosh, 2018). Testing configuration, however, was observed by Rzig et al. (2022) to be prevalent across a large number of projects. Conversely, Gallaba and McIntosh (2018) highlighted that, while deployment configuration is less common, it can demand considerable configuration code when implemented. Though we expect testing and deployment to have more adoption and complexity in Android apps, due to the variety of platforms requiring testing and the complexities of app store deployments, our findings suggest otherwise.
@@
Figure 4. Boxplots showing the total number of directives across CI/CD services, alongside the respective ratios of directives for each category. Icons:Refer to caption GitHub Actions  •  Refer to caption Travis CI  •  Refer to caption CircleCI  •  Refer to caption GitLab CI/CD
Observation 1.3: Nearly half of the projects implement automated testing processes, with the predominant standard testing tool, through Gradle, accounting for about 78% of the total. Among the analyzed projects, 1,223 (47.7%) use automated testing processes as part of the CI/CD pipeline. Table 4 shows the top 10 testing tools used by the projects in our dataset. We should note that multiple testing tools can be used in a single project, resulting in cumulative percentages that exceed 100%. The results show a preference for adopting testing practices through standard tools like Gradle in Android projects. The results are consistent across CI/CD services. This trend likely arises from the ease of integration and familiarity, leading to more efficient testing workflows. We observe that unit testing is by far the most commonly used practice, with 78.3% of projects incorporating it. This highlights the importance of unit testing in software development, as it enables developers to verify code correctness at a granular level. This suggests that unit testing is deemed essential for maintaining code reliability and is likely prioritized due to its effectiveness in identifying defects early and its seamless integration with continuous integration CI/CD pipelines.
Android-specific tools, such as Android Lint (28.7%), testing on connected devices (14.5%), and the Android Emulator (9.6%), show a relatively lower adoption compared to unit testing, despite their role in ensuring platform compliance and optimizing performance. The greater reliance on Android Lint suggests an interest of Android developers in performing static code analysis to detect issues before runtime. However, the low adoption rates of connected devices and emulators might indicate that many teams prefer more lightweight and automated approaches, possibly relying on simulated environments or static checks before moving to the next steps.
@@
Finally, tools like Firebase [^22] https://firebase.google.com (5.1%), Fastlane [^23] https://fastlane.tools (2.2%), and SonarQube [^24] https://www.sonarsource.com (2.1%) show even lower adoption rates, despite their popularity as third-party testing platforms. Firebase’s usage likely reflects a focus on cloud-based testing and monitoring, particularly for mobile apps, but its usage remains relatively limited in our datasets. Fastlane and SonarQube, which facilitate CI/CD automation and comprehensive code quality analysis, respectively, appear to be more specialized tools that are leveraged in specific cases, possibly due to their integration complexity.
@@
Testing Tool Configuration Examples # projects (%)
@@
Compare & Contrast with Prior Research Findings (1.3): Similar to our findings, Zhang et al. (2022) indicated that Gradle is a prevalent testing tool. However, Maven, which was observed to be dominant in Java projects (Zhang et al., 2022), appears to be less common for Android apps. In addition, Rzig et al. (2022) reported various tools for testing machine learning projects developed with Python, underscoring the need to investigate Android-specific testing tools. We discover several tools unique to Android testing (e.g., Android Lint and Android Emulator) that are not commonly used in regular software testing.